# Post-detection inference for sequential changepoint localization

### v1

@@ 5 Experiments
We now experimentally validate our algorithms for constructing confidence sets in order to evaluate their performance across various settings and demonstrate their practical utility.
@@ 5.1.1 Known pre- and post-change (Setting I)
We consider both Normal and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using the CUSUM change detection method defined in (B.1) with thresholds at $A=1000,10000$ (which controls ARL at level $A$ (Tartakovsky et al.,, 2014)). After detection, we employ Algorithm 2 to construct confidence sets for $T$. Figures 5.1 and 5.2 provide visualizations of the confidence sets and point estimates (2.7) of the changepoints across $5$ runs.
Figure 5.1: Setting I (known pre- and post-change Gaussian distributions): The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$ for $T=100$ (top row) and $T=500$ (bottom row). Change is detected using the CUSUM detector (B.1) with $A=10000$ (top row) or $A=1000$ (bottom row). The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
Figure 5.2: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois$(1)$ and the rest from Pois$(2)$, with $T=100$ (top row) or $T=500$ (bottom row). Change is detected using CUSUM detector (B.1) with $A=10000$ (top row) or $A=1000$ (bottom row). The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.3.
Table 5.3: Setting I: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (B.1) with threshold $A=1000$, while varying $T$.
$T$ $F_{0}$ $F_{1}$
@@
100 $N(0,1)$ $N(1,1)$ 0.96 0.94 12.34 2.85 13.97
100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.95 17.45 3.98 15.72
500 $N(0,1)$ $N(1,1)$ 0.95 0.86 12.57 2.62 13.22
500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.83 18.72 3.96 16.14
@@ 5.1.2 Known pre-change and composite post-change (Setting II)
We perform experiments for Gaussian mean-change scenarios having pre-change distribution as $N(0,1)$, and the true (unknown) post-change data distribution as $N(1,1)$, with the changepoint at $T=100,500.$ For detection, we employ a weighted CUSUM detector (3.16) with thresholds at $A=1000,10000$ (which controls ARL at $A$). For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.75\}$ and $\theta=0.9$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we construct confidence sets (as defined in (3.11)) for $T$. Fig. 5.4 provides visualizations of the confidence sets (3.11) and point estimates (3.5) of the changepoint across $5$ random runs.
@@
We report the average size and the conditional and unconditional coverage across $100$ independent runs in (5.5). For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.3). We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
Table 5.5: Setting II: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.11) with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ (3.5) and average detection delay (given no false detection) of the change detection method used (3.16) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and composite post-change model.
@@
100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 23.25 3.95 16.87
100 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 20.11 3.67 16.21
500 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.87 30.83 3.45 17.63
500 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.86 28.56 3.48 15.81
@@ 5.1.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, with the changepoint at $T=100$ and $T=500$. For change detection, we employ a weighted CUSUM-type detector (defined in (4.17)), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in setting II is considered. We set the thresholds at $A=1000,10000$. This is a special case of “e-detector” and is known to control ARL at level $A$ (Shin et al.,, 2023).
After detection, we construct confidence sets (as defined in (4.12)) for $T$ (assuming $T\neq 1$). We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs in Table 5.6.
Table 5.6: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ (4.6) and average detection delay (given no false detection) of the change detection method used (4.17) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and pre- and post-change model.
@@
100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 25.51 4.36 25.81
100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 22.97 4.19 23.15
500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.95 0.95 30.74 4.03 24.48
500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 29.16 4.07 23.07
However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.3). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table 5.3), exhibiting both higher coverage and larger size.
Fig. 5.7 visualizes the confidence sets (4.12) and point estimates (4.6) of the changepoint across $5$ random runs.
@@ 5.2.1 Known pre- and post-change (Setting A)
with thresholds at $A=1000$. Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality (Ville,, 1939), the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we employ Algorithm 1 to construct confidence sets for $T$. We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.8. We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
Table 5.8: Setting A: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 1 with $B=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (defined in (5.1)) with $A=1000$, while varying the changepoint $T$.
@@ 5.2.2 Known pre- and unknown post-change (Setting B)
with thresholds at $A=1000$. The weights are chosen just as described in Section 5.1.2. Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct confidence sets for $T$ as defined in (3.12). We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.9. We also report the average absolute deviation of $\hat{T}$ (defined in (3.5)) and delay (given no false detection) in the table.
Table 5.9: Setting B: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.12) with $B=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation of $\hat{T}$ (defined in (3.5)) and average detection delay (given no false detection) of the change detector used (as defined in (5.2)) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and composite post-change model.
@@ 5.2.3 Composite pre- and post-change (Setting C)
We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) Larsson et al., (2025); Grünwald et al., (2024); Lardy et al., (2024) being the representative of the pre-change class) defined as
@@
with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$. This is a stop time of an ”e-process” (Ramdas et al.,, 2023) and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. The weights are chosen just as described in Section 5.1.3. After detection, we construct confidence sets for $T$ as defined in (4.13) (assuming $T\neq 1$). We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.10. We also report the average absolute deviation of $\hat{T}$ (defined in (4.6)) and delay (given no false detection) in the table.
Table 5.10: Setting C: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $B=100$ and $\alpha=0.05,\beta=\gamma=0.025,L=\infty$, the average absolute deviation of $\hat{T}$ (defined in (4.6)) and average detection delay (given no false detection) of the change detector used (as defined in (5.3)) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and pre- and post-change models.
@@ 6 Conclusion
We developed a framework to construct a confidence set for the unknown changepoint $T$, providing conditional coverage guarantees (conditioned on the event that the stopping occurs after the changepoint) when the pre- and post-change distributions are known. We also extended our framework to handle composite pre- and post-change distributions and provide conditional coverage guarantees, demonstrating its good empirical performance. We impose no specific conditions on the change detection algorithms; they can function as entirely black-box or heuristic methods; we only need to be able to run the algorithm on simulated data sequences. By filling this significant gap in post-detection analysis, our work provides crucial tools for precise and timely changepoint localization in real-time data streams. Our framework offers important implications for applications where rapid and accurate inference is critical, paving the way for more effective changepoint localization in sequential settings. While our method is designed for parametric pre- and post-change models in the partitioned setting, future work can explore its extensions to the non-partitioned setting and nonparametric models.
@@ A.1 Omitted proofs from Section 2
Hence, $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})\iff M_{t}$ is one among the $\lceil(1-\alpha)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$. So, for any fixed $\alpha\in(0,1)$,
$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-\alpha)(B+1)\rceil}{B+1}\geq 1-\alpha,\text{ for all }t.$ (A.1)
@@
Note that $\mathcal{C}^{*}=\{1\leq t\leq\tau:M_{t}\leq C_{t,L}\}\cup\{\tau+1,\tau+2,\cdots\}=\{t\geq 1:M_{t}\leq C_{t,L}\}$, since we defined $M_{t}=-\infty\leq C_{t,L}$, for all $t>\tau.$
@@
$\mathbb{P}_{T}(T\in\mathcal{C}^{*})=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})=\mathbb{P}_{T}(M_{T}\leq C_{t,L})\geq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})\geq 1-\alpha,$ (A.2)

### v2

@@ 5 Experiments
We validate our algorithms across various settings and demonstrate their practicality. Experiments below are for Gaussians; Poisson is handled in Appendix B.
@@ 5.1.1 Known pre- and post-change (Setting I)
We consider Gaussian mean-change from $N(0,1)$ (pre-change) to $N(1,1)$ (post-change) having the true (unknown) changepoint at $T=100,500$. We employ the CUSUM change detection method (C.1) with thresholds at $A=1000,10000$ (which controls ARL at level $A$ [23]). After detection, we employ Algorithm 2 to construct confidence sets for $T$, with $B=N=100$, $\alpha=0.05$. Fig. 5.1 visualizes the confidence sets and point estimates (2.7) across $5$ runs. Table 5.2 has the average size and the coverages across $100$ independent runs.
Figure 5.1: Setting I (known pre- and post-change Gaussian distributions): The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$ for $T=100$ (top row) and $T=500$ (bottom row). Change is detected using the CUSUM detector (C.1) with $A=10000$ (top row) or $A=1000$ (bottom row). The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
Table 5.2: Known pre- and post-change (Setting I, finite ARL)
$T$ A
@@
100 10000 0.96 0.96 12.21 2.97 17.63
100 1000 0.96 0.94 12.34 2.85 13.97
500 10000 0.95 0.95 12.88 2.59 17.26
500 1000 0.95 0.90 12.57 2.62 13.22
@@ 5.1.2 Known pre-change and composite post-change (Setting II)
We perform experiments for Gaussian mean-change scenarios having pre-change distribution as $N(0,1)$, and the true (unknown) post-change data distribution as $N(1,1)$, with the changepoint at $T=100,500.$ For detection, we employ a weighted CUSUM detector (defined in (3.15)) with thresholds at $A=1000$ (which is known to control ARL at level $A$). For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.75\}$ and $\theta=0.9$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we construct confidence sets (as defined in (3.11)) for $T$. We report the average size and coverages across $100$ independent runs in Table 5.4. Fig. 5.3 provides visualizations of the confidence sets (3.11) and point estimates (3.5) of the changepoint across $5$ random runs.
@@
Table 5.4: Known pre-change but unknown post-change (Setting II, finite ARL)
@@
100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 26.25 3.95 16.87
100 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 24.11 3.67 16.21
500 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.89 30.83 3.45 17.63
500 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.91 28.56 3.48 15.81
For $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in much lower unconditional coverage. However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.2). We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
@@ 5.1.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, with the changepoint at $T=100$ and $T=500$. For change detection, we employ a weighted CUSUM-type detector (4.17), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in setting II is considered. We set the thresholds at $A=1000,10000$. This is a special case of “e-detector” and is known to control ARL at level $A$ [20]. After detection, we construct confidence sets (4.12) for $T$ (assuming $T\neq 1$), with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$. Fig. 5.5 visualizes the confidence sets (4.12) and point estimates (4.6) of the changepoint across five random runs.
Table 5.6 contains the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs. The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.2). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table 5.2), exhibiting both higher coverage and larger size.
Table 5.6: Composite pre- and post-change (Setting III, finite ARL)
@@
100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 27.84 4.36 25.81
100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 25.71 4.19 23.15
500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 31.45 4.03 24.48
500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 30.66 4.07 23.07
@@ 5.2.1 Known pre- and post-change (Setting A)
with thresholds at $A=1000$. Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [24], the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we employ Algorithm 1 to construct confidence sets for $T$, with $B=100$ and $\alpha=0.05$. We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.7. We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
Table 5.7: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
@@ 5.2.2 Known pre- and unknown post-change (Setting B)
with thresholds at $A=1000$. The weights are chosen just as described in Section 5.1.2. Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct confidence sets for $T$ as defined in (3.12). We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.8, with $B=100$ and $\alpha=0.05,\beta=0.025,L=\infty$. We also report the average absolute deviation of $\hat{T}$ in (3.5) and delay (given no false detection) in the table.
Table 5.8: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
@@ 5.2.3 Composite pre- and post-change (Setting C)
We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) [13, 6, 12] being the representative of the pre-change class) defined as
@@
with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$. This is a stop time of an ”e-process” [17] and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. The weights are chosen just as described in Section 5.1.3. After detection, we construct confidence sets for $T$ as defined in (4.13) (assuming $T\neq 1$). The pre- and post-change data-generating distributions are $N(0,1)$ and $N(0,1)$ respectively, but neither of which is assumed known, and we vary the pre- and post-change classes. We use $B=100$ and $\alpha=0.05,\beta=\gamma=0.025,L=\infty$. We report the average size and coverage rates across $100$ independent runs in Table 5.9. We also report the average absolute deviation of $\hat{T}$ in (4.6) and delay (given no false detection). The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
Table 5.9: Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
@@ 6 Conclusion
We developed a framework to construct a confidence set for the unknown changepoint $T$ after a sequential detection algorithm declares a change at a stopping time, providing conditional coverage guarantees (conditioned on the event that the stopping occurs after the changepoint) when the pre- and post-change distributions are known. We also extended our framework to handle composite pre- and post-change distributions and provide conditional coverage guarantees, demonstrating its good empirical performance. We impose no specific conditions on the change detection algorithms; they can function as entirely black-box or heuristic methods; we only need to be able to run the algorithm on simulated data sequences. By filling this significant gap in post-detection analysis, our work provides crucial tools for precise and timely changepoint localization in real-time data streams.
@@ A.1 Omitted proofs from Section 2
[…]
$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-\alpha)(B+1)\rceil}{B+1}\geq 1-\alpha.$ (A.1)
@@
Note that $\mathcal{C}^{*}=\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\cup\{\tau+1,\tau+2,\cdots\}=\{t\geq 1:M_{t}\leq C_{t,L}\}$, since we defined $M_{t}=-\infty\leq C_{t,L}$, for $t>\tau.$
@@
$\mathbb{P}_{T}(T\in\mathcal{C}^{*})=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})=\mathbb{P}_{T}(M_{T}\leq C_{t,L})\geq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})\geq 1-\alpha,$